Database last updated on Tue, 27 Jan 2026 17:40:45 +0000 by localhost
Critical warnings from b-files' comments for days between 2025112 and 2025112 (0 sec, 7 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| hg: | Lamp not on... hg test terminated F1= 0 F4= 326 | 1 | 1 | B11225.254 | TORON | #254 | 00:32:05 | | 1 | -1 |
| hg: | Lamp not detected f1= 0 f4= 326 | 1 | 1 | B11225.254 | TORON | #254 | 00:32:06 | | 1 | -1 |
| rl: | 2025 111 05 42 47 New value set for watermark SUPPLY.VOLTAGE.HIGH | 14 | 3 | B11225.254 | TORON | #254 | 05:28:12 | | 1 | -1 |
| rl: | 2025 111 05 55 19 New value set for watermark LOOP.RUN.TIME: 655.000019 millisec. | 10 | 3 | B11225.254 | TORON | #254 | 05:28:17 | | 1 | -1 |
| rl: | 2025 112 11 52 56 PMT counter failure | 1 | 1 | B11225.254 | TORON | #254 | 05:28:43 | | 1 | -1 |
| rl: | 2025 113 00 19 38 Reset No -> Warm from 2025 113 00 19 34 until 2025 113 00 19 35 . Counts: 0,1 | 5 | 3 | B11225.254 | TORON | #254 | 05:28:49 | | 1 | -1 |
| ds: | Suppressed oscillation of filter wheel. | 6 | 5 | B11225.254 | TORON | #254 | 13:29:26 | | 1 | -1 |
Instrument constant changes for the week before 2025112 (2 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 2 rows)
Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)
| Serial | YYYYJJJ | HG large | max correction |
| 254 | 2025112 | 2 | 16.7 |
Non-critical warnings from b-files' comments for days between 2025112 and 2025112 (0 sec, 4 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 111 05 41 49 Reset Warm -> Tepid from 2025 111 05 41 43 until 2025 111 05 41 44 . Counts: 1,0 | 1 | 3 | B11225.254 | TORON | #254 | 05:28:09 | | 1 | -1 |
| rl: | 2025 112 06 21 00 New value set for watermark LOOP.RUN.TIME: 715.000010 millisec. | 3 | 3 | B11225.254 | TORON | #254 | 05:28:38 | | 1 | -1 |
| rl: | 2025 112 06 21 00 New value set for watermark LOOP.RUN.TIME: 715.000010 millisec. | 3 | 3 | B11225.254 | TORON | #254 | 05:28:38 | | 1 | -1 |
| rl: | 2025 113 00 20 59 New value set for watermark SUPPLY.VOLTAGE.HIGH | 2 | 3 | B11225.254 | TORON | #254 | 05:29:05 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | ds | dt | ed | fz | hg | rs | sl | uv | zs |
| 112 | TORONTO | #254 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.281) | 70 | 2 | 2 | 5 | 22 | 2 | 7 | 29 | 40 |
(0 sec)
Number of days since last dome/window cleaning (0 sec, 1 rows)
Missing files (0 sec, 1 rows)
| Serial | last data | days ago |
| 254 | 2025113 | 0-1 |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 254 | 15 | 0 | 1 | 0 | 0 | 0 | 0 | 0 | 0 |
Standard lamp intensity analysis (0 sec, 1 rows)
| Serial | min | max | filter | fixable | 30-day history |
| #254 | 2552690 | 2554350 | 0 | yes |